# Transferring disentangled representations: bridging the gap between synthetic and real images

### v1

@@ 1 Introduction
Developing meaningful, reusable and efficient representations is a critical step in representation learning [1, 56, 55, 64]. Disentangled Representation Learning (DRL) [1, 39, 24, 64] aims to learn models that can identify and disentangle underlying Factors of Variation (FoVs), hidden in the observable data, encoding them in an interpretable and compact shape [31, 9, 1, 69], independently from the task at hand [22, 38, 63, 64]. Moreover, DRL enhances explainability, robustness, and generalization capacity across various applications [64]. Disentangled representations have been shown useful for various downstream tasks, such as FoVs prediction [40, 39], image generation [68, 46, 42, 41, 57] and translation [21, 19, 37], fair classification [54, 38], abstract reasoning [63, 60], domain adaptation [35], and out-of-distribution (OOD) generalization [11, 20].
@@
In this work, we propose the adoption of Disentangled Representation (DR) transfer to deal with complex realistic/real dataset. DL tranferring was explored in [20], where Source models learnt in an unsupervised manner were transferred to a Target dataset, by transferring hyperparameters. The authors observed a limited effectiveness in the direct transfer of representations. Instead, Dittadi et al. [11] found out that disentangled representation can help in OOD generalization from a simulated to a smaller real dataset. In both cases, the study involved very specific types of dataset, built to emulate the real one in every detail. Recently, Fumero et al. [18] addressed disentanglement in real data without the need for FoVs annotation, leveraging the knowledge extracted from a diversified set of supervised tasks to learn a common disentangled representation to be transferred to real settings. We follow a different direction, setting up a very straightforward and generalizable procedure: we resort to a weakly supervised approach [40, 25, 26] to learn DRs on Source datasets where the FoVs are known and annotated, to then transfer (with no supervision) such representation to a Target dataset where the FoVs are not known or available. Our final aim is to consider real datasets as a Target, while synthetic data (where FoVs annotation is easy to obtain) can be employed as a Source.
@@ 2.2 Our metric: OMES
OMES (Overlap Multiple Encoding Scores) is an intervention-based metric measuring the quality of factor encoding in the representation while providing information about its structure: we measure modularity, analyzing how the FoVs overlap, and compactness, detecting and quantifying how a factor is encoded in the dimensions of the representation. To the best of our knowledge, the only metrics capturing more than one property are DCI [13] and the very recent InfoMEC [25]. Differently from DCI, our metric is intervention-based with no influence on the choice of the specific classifier that may inevitably impact the results, as observed in [7]. With respect to InfoMEC, that must be applied to quantized latent codes, our metric is more general and accepts continuous latents.
Algorithm 1 Compute association matrix between dimensions and FoVs
@@
Given an image $X$, $\Phi$ is its mapping into a $d-$dim. latent disentangled space, $\Phi(X)=r$, $r\in\mathbb{R}^{d}$. We discard dimensions whose empirical standard deviation is extremely small ($<0.05$), meaning that the dimensions are inactive [65, 10].This leaves us with a subset of $m\leq d$ active dimensions, to which we will refer in the following.
@@
Let $D$ be a dataset formed by image pairs, $D=\{(X_{i}^{1},X_{i}^{2},k_{i})\}_{i=1}^{N}$, where $X_{i}^{1},X_{i}^{2}$ are two images that differ for only the FoV $k_{i}$ (e.g. object color).
OMES requires computing a weighted association matrix $S$ between the dimensions of the representation and the FoVs, with higher association values if the factor is encoded in a certain dimension. The procedure is described in Algorithm 1. We consider the representations of the image pairs in dataset $D$, obtaining $D_{\Phi}$. In matrix notation we may write it as $D_{\Phi}=[\boldsymbol{R}^{1},\boldsymbol{R}^{2},\boldsymbol{k}]$, where the pair $\boldsymbol{R}^{1}$ and $\boldsymbol{R}^{2}$ are $N\times m$-dimensional matrices with each row $i$ is the representation of the $i$-th image pair $\Phi(X^{1}_{i})=r_{i}^{1}$ and $\Phi(X^{2}_{i})=r_{i}^{2}$ respectively. For each FoV $k$ we extract the rows of $D_{\Phi}$ such that the $i$-th entry of vector $\boldsymbol{k}$ is $k_{i}=k$, we call this set $D_{\Phi}^{k}$.
Each matrix entry $S[h,j]$ relates a dimension $h$ of the estimated disentangled representation with a FoV $j$. Its value is in the range $[0,1]$ with elements close to $1$ corresponding to a dimension that effectively captures the variations of a FoV. The association is based on a correlation analysis: since the samples from $D_{\Phi}^{k}$ are paired to differ only for the FoV $k$, we expect a good representation not to correlate where such FoV is encoded. To ease interpretability, we transform the obtained values (see Algorithm 1, line 6) so that high values denote a strong association between dimension and FoV.
When the model exhibits perfect disentanglement, each row and column of the association matrix $S$ present just one element with a high association, corresponding to the only dimension where the factor is encoded. We thus measure the level of disentanglement through similarity with an ideal array, where the association matrix shows all 0s but in the positions of the correct associations, where there are 1s. We rely on the above considerations to derive our metric as a linear combination of two main contributions. The Overlap Score (OS) penalizes the overlap of different FoVs in the same dimensions (Algorithm 2 — in this case, each row of $S$, associated with a dimension, is compared with the ideal array) and hence measures Modularity, while the Multiple Encoding Score (MES) penalizes the encoding of the same factor into different dimensions (Algorithm 3 — in this case each column of $S$ corresponding to a FoV is compared with the ideal array) measuring Compactness. In both algorithms, we derive a vector summarizing the contribution of all dimensions for the considered FoV. The final score in $[0,1]$ (higher values meaning higher disentanglement) can be obtained with a pooling (either MAX or AVERAGE) on the vector. The OMES metric is computed as
@@ 2.3 OMES assessment
We now analyze OMES, extending previous studies on the unsupervised [39] and weakly supervised [40] setting. As for the unsupervised case, we exploit available 5400 trained models from [39] (3 datasets, 6 values for $\beta$, 50 random seeds, 6 unsupervised methods:$\beta$-VAE [24], FactorVAE [27], $\beta$-TCVAE [8], DIP-VAE-I [32], DIP-VAE-II[32], AnnealedVAE [5]); in this section we report an analysis on Noisy-dSprites, the remaining 2 benchmarks can be found in the Appendix B.1 and B.2. Instead, for the weakly supervised case trained models are not available, so we reproduce the models as in [40] training them on Shapes3D and on others that can be found in Appendix B.3.
@@
Fig. 1 (Left) shows the metric scores for different values of $\beta$ keeping the different FoV separated: the FoVs less affected by reconstruction (e.g. PosX and PosY) exhibit an increasing disentanglement score as $\beta$ grows. On the other hand, Shape and Orientation present a maximum value around $\beta=6$ and then decrease because they are more susceptible to the reconstruction quality, which degrades for larger values of $\beta$. In Fig. 1 (Center Left) an association matrix S is generated from one of the unsupervised models, $\beta$-VAE, trained with $\beta=6$: Shape and Orientation are encoded in the same dimension (overlapping) and produce lower values because of the reconstruction, while PosX and PosY are encoded in multiple dimensions and mostly overlapping. Scale does not seem to be well represented. In Fig. 1 (Center Right) we report the association matrix obtained by a weakly-supervised model (Ada-GVAE): the factors PosX, PosY and Scale are disentangled while Shape and Orientation are encoded in the representation with high intensity in different dimensions, with overlaps. Fig.1 (Right) shows OMES values ($\alpha=0.5$) computed over synthetic association matrices $S$, obtained by perturbing the ideal (diagonal) one. The perturbation aims to simulate 3 scenarios: noisy matrices where the disentanglement can be more or less strongly derived; models exhibiting partial disentanglement; models with no disentanglement. As it can be appreciated, the metric score nicely reflects the disentanglement intensity.
@@
Agreeement with performance metrics. Similarly to [40], we consider the more informative weakly-supervised setting and discuss the rank correlation between our metric and performance evaluations (ELBO, reconstruction loss, and test error of FoVs classifier). Our analysis, reported in Fig. 2 (right) shows that OMES performs similarly to DCI, negatively correlated with the Reconstruction loss, and positively with the ELBO. It also correlates with the performances of GBT10000 (the classifier we will use in the experiments) while it mildly does with MLP10000. This empirical evidence is in line with what was observed in [11]. The correlation with the classification score is a sign that OMES is able to capture the property of expliciteness of the representation, although it is not directly measured by our metric. It is worth mentioning the correlation of OMES with the performance metrics is more stable than what is obtained by other metrics across different datasets (see the Appendix B.3).
@@ 3 Transferring disentangled representations
Fully unsupervised disentangled representation learning has been shown unsatisfactory in many scenarios [39]. However, annotating the FoVs can be a very critical and uncertain process. In this section, we propose a general-purpose methodology for transferring disentangled representations learned from supervised synthetic or simulated data to an unsupervised dataset (in terms of the FoVs). This approach allows us to evaluate the effectiveness of disentangled representations transfer, and its potential in real world applications.
@@ 3.2 Datasets
Coil is derived from Coil100 [45]. The original dataset contains 7200 real color images of 100 objects. The objects were placed on a motorized turntable against a black background. The turntable was rotated to vary object pose w.r.t. a fixed camera, producing self-occlusions and 2D silhouette changes. We augment the original dataset with two additional FoV, a planar rotation (9 angles) and a scaling (18 values). Therefore, we identify 4 FoV (Objects, Pose, Rotation and Scale) that, by construction, are independent. To consider in our analysis a real dataset visually related to dSprite, we derived a binary version, called Coil(bin), by applying Otsu’s thresholding [47] to each image of Coil.
@@ 3.3 Experimental analysis
The column Pruned highlights the two different representation modalities: if the classifier is trained on the whole representation (), or using only one dimension, the one showing the strongest encoding of a certain FoV according to the OMES metric (). As already mentioned, a good performance of the former is an indication of explicitness, while the latter is a positive sign of compactness. Tables also report metrics assessing Modularity (our MES and DCI) and Compactness (our OS and MIG).
Table 2: Target dataset: dSprites and variant (see text).
@@
(1) Synthetic to synthetic. As a baseline, we consider the case in which both Source and Target datasets are synthetic and we have access to the annotation of the FoVs, they are DRL-compliant. If Source and Target have the same FoVs (S=dSprites with T=Noisy-dSprites or S=Color-dSprites with T=Noisy-Color-dSprites, see Table 2) we observe the following: pruning the representation to just one dimension maintains, on average, stable performances, showing that the compactness of the representation is preserved for the Target dataset, both before and after fine-tuning. Fine-tuning allows for improved performance in terms of explicitness preserving the remaining properties of the representation, also in the case of the pruned representation. The Orientation FoV is difficult in these datasets as it suffers from reconstruction errors. We increase complexity by adding a new FoV to the Target dataset (S=dSprite with T=Color-dSprite, see Table 2). All FoVs in common between Source and Target are effectively classified, again except Orientation. As for the new FoV (Color), we report lower performances, but we can appreciate a significant improvement with fine-tuning if we exploit a global representation. Instead, we observe a lower improvement with the pruned representation, suggesting that the new factor is not encoded in one single dimension.
@@
Table 3: Target dataset: Shapes3D (see text).
@@
Discussion. Disentanglement transfers well between synthetic datasets with the same FoVs, w.r.t. all the properties. If the Target includes new FoVs, fine-tuning is necessary for the new FoV, but also for the entire representation, as compactness and modularity are partially degraded by the new FoV. When the Source and Target become significantly different, fine-tuning is also beneficial. We can conclude that when both Source and Target are synthetic and DRL-compliant, the properties of disentangled representation are preserved before and after finetuning.
@@
With the same Target, we assess a representation which is incorporating some level of 3D information ( S=Shapes3D, with T=Coil, Table 4). This choice does not bring any benefit since the synthetic 3DShape includes a very simplified form of pose variation. Orientation accuracy degrades, as it is not captured by the Source (indeed, with fine tuning this performance improves). Considering the Target dataset, it clearly presents several new challenges w.r.t. Source ones, we produce a last experiment with a simplified binary version of the Target, meant to be more similar to the binary images in dSprite. In this case, some FoVs are very well represented (Orientation and Scale), while Object presents low performances due to the decrease in the image descriptive power caused by binarization.
We now consider another real Target, RGBD-Object(see Table 5). As for the former, we consider Color-dSprite as a synthetic source: here, the same considerations about the Pose we discussed before are valid, and on the other FoVs, we observe the global representation is effective (and marginally improved by fine-tuning), while the pruned representation leads to lower performances, as a sign the representation is not perfectly disentangled on the Target. Here again, we investigate the possibility of transferring to simplified versions of the dataset (in this case, we consider the Depth channel only). The FoV Elevation improves significantly as it is not directly affected by color information. Concerning Modularity (Table 4 and Table 5), fine-tuning seems to have a small influence. In particular we observe a small degradation for RGBD-Objects (Table 5), coherent with the challenges of a real dataset with unknown/hidden factors perturbing the encoding of the FoV.
Table 4: Target dataset: Coil100 and variants (see text).
@@
Table 5: Target dataset: RGBD-Objects and variants (see text).
@@
(3) Real to Real. We conclude by discussing the possibility of transferring from a DRL-compliant real dataset to another real one. As a first task, we consider as a Source a simplified version of the Target (specifically, S= Coil-binary, with T=Coil): the source should encode the factors not related to RGB, while the finetuning should improve the disentanglement and the explicitness of the representation. However, this is not the case with Coil100, whose representations degrade the Modularity, and the finetuning only affects the entire representation. We then consider a larger variation between Real Source and Real Target (specifically, S=Coil with T=RGBD-Object, see Table 5): we obtain similar results to those of Color-dSprites as Source Dataset (comparable Explicitness), with a reduction on the performances obtained by the pruned representation. Notice that adopting the binary Coil as a Source causes only a limited reduction in Explicitness, and this was somewhat unexpected as we have a large gap in complexity between Source and Target. Our experiments did not consider RGBD-Objects acting as a Source dataset, not being DRL-compliant.

### v3

@@ 1 Introduction
Developing meaningful, reusable and efficient representations is a critical step in representation learning [1, 58, 57, 67]. Disentangled Representation Learning (DRL) [1, 40, 24, 67] aims to learn models that can identify and disentangle underlying Factors of Variation (FoVs), hidden in the observable data. These models encode them in an interpretable and compact shape [31, 9, 1, 73], independently from the task at hand [22, 39, 66, 67]. Moreover, DRL enhances explainability, robustness, and generalization capacity across various applications [67]. Disentangled representations have been shown useful for various downstream tasks, such as FoVs prediction [41, 40], image generation [72, 48, 43, 42, 59] and translation [21, 19, 38], fair classification [56, 39], abstract reasoning [66, 63], domain adaptation [35], and out-of-distribution (OOD) generalization [11, 20].
@@
In this work, we propose the adoption of Disentangled Representation (DR) transfer to deal with complex realistic/real dataset. DL transferring was explored in [20], where Source models learnt in an unsupervised manner were transferred to a Target dataset, by transferring hyperparameters. The authors observed a limited effectiveness in the direct transfer of representations. Instead, Dittadi et al. [11] found out that disentangled representation can help in OOD generalization from a simulated to a smaller real dataset. In both cases, the study involved very specific types of dataset, built to emulate the real one in every detail. Recently, Fumero et al. [18] addressed disentanglement in real data without the need for FoVs annotation, leveraging the knowledge extracted from a diversified set of supervised tasks to learn a common disentangled representation to be transferred to real settings. We follow a different direction, setting up a very straightforward and generalizable procedure: we resort to a weakly supervised approach [41, 25, 26] to learn DRs on Source datasets where the FoVs are known and annotated, to then transfer (with no supervision) such representation to a Target dataset where the FoVs are not known or available. Our final aim is to consider real datasets as a Target, while synthetic data (where FoVs annotation is easy to obtain) can be employed as a Source.
@@ 2.2 Our metric: OMES
OMES (Overlap Multiple Encoding Scores) is an intervention-based metric measuring the quality of factor encoding in the representation while providing information about its structure: we measure modularity, analyzing how the FoVs overlap, and compactness, detecting and quantifying how a factor is encoded in the dimensions of the representation.
Algorithm 1 Compute association matrix $S$ between dimensions and FoVs
@@
Given an image $X$, with $\Phi$ its mapping into a $d-$dimensional latent disentangled space, $\Phi(X)=r$, $r\in\mathbb{R}^{d}$. We discard dimensions whose empirical standard deviation is extremely small ($<0.05$), meaning that the dimensions are inactive [68, 10].This leaves us with a subset of $m\leq d$ active dimensions, to which we will refer in the following.
@@
Let $D$ be a dataset formed by image pairs, $D=\{(X_{i}^{1},X_{i}^{2},k_{i})\}_{i=1}^{N}$, where $X_{i}^{1},X_{i}^{2}$ are two images that differ for only the FoV $k_{i}$.
OMES requires computing a weighted association matrix $S$ between the dimensions of the representation and the FoVs, with higher association values if the factor is encoded in a certain dimension (see Algorithm 1): we consider the representations of the image pairs in dataset $D$, obtaining $D_{\Phi}$. In matrix notation we may write it as $D_{\Phi}=[\boldsymbol{R}^{1},\boldsymbol{R}^{2},\boldsymbol{k}]$, where the pair $\boldsymbol{R}^{1}$ and $\boldsymbol{R}^{2}$ are $N\times m$-dimensional matrices with each row $i$ is the representation of the $i$-th image pair $\Phi(X^{1}_{i})=r_{i}^{1}$ and $\Phi(X^{2}_{i})=r_{i}^{2}$ respectively. For each FoV $k$ we extract the rows of $D_{\Phi}$ such that the $i$-th entry of vector $\boldsymbol{k}$ is $k_{i}=k$, we call this set $D_{\Phi}^{k}$. Each entry $S[h,j]$ of the association matrix relates a dimension $h$ of the estimated disentangled representation with a FoV $j$. Its value is in the range $[0,1]$ with elements close to $1$ corresponding to a dimension that effectively captures the variations of a FoV. The association is based on a correlation analysis: since the samples from $D_{\Phi}^{k}$ are paired to differ only for the FoV $k$, we expect a good representation not to correlate where such FoV is encoded. To ease interpretability, we transform the obtained values (see Algorithm 1, line 6) so that high values denote a strong association between dimension and FoV.
When the model exhibits perfect disentanglement, each row and column of the association matrix $S$ present just one element with a high association, corresponding to the only dimension where the factor is encoded. We thus measure the level of disentanglement through similarity with an ideal array, where the association matrix shows all 0s but in the positions of the correct associations, where there are 1s.
We rely on the above considerations to derive our metric as a linear combination of two main contributions. The Overlap Score (OS) penalizes the overlap of different FoVs in the same dimensions (Algorithm 2 — in this case, each row of $S$, associated with a dimension, is compared with the ideal array) and hence measures Modularity, while the Multiple Encoding Score (MES) penalizes the encoding of the same factor into different dimensions (Algorithm 3 — in this case each column of $S$ corresponding to a FoV is compared with the ideal array) measuring Compactness. In both algorithms, we derive a vector summarizing the contribution of all dimensions for the FoV.
The final score in $[0,1]$ (higher values meaning higher disentanglement) can be obtained with a pooling (either MAX or AVERAGE) on the vector. The OMES metric is computed as
@@
With $\alpha=0$, OMES only measures the Compactness of the representation (MES component); with $\alpha=1$, instead, our metric measures the Modularity only (with OS). Values of $\alpha$ in the interval $(0,1)$ can be used to balance the importance of both contributions.
Relation with existing metrics. To the best of our knowledge, the only metrics capturing more than one property are DCI [13] and the very recent InfoMEC [25]. Differently from DCI, our metric is intervention-based with no influence on the choice of the specific classifier that may inevitably impact the results, as observed in [7]. With respect to InfoMEC, that must be applied to quantized latent codes, our metric is more general and accepts continuous latents.
@@ 2.3 OMES assessment
We now analyze OMES, extending previous studies on the unsupervised [40] and weakly supervised [41] setting. As for the unsupervised case, we exploit available 5400 trained models from [40] (3 datasets, 6 values for $\beta$, 50 random seeds, 6 unsupervised methods:$\beta$-VAE [24], FactorVAE [27], $\beta$-TCVAE [8], DIP-VAE-I [32], DIP-VAE-II[32], AnnealedVAE [5]); in this section we report an analysis on Noisy-dSprites, the remaining 2 benchmarks can be found in the Appendix B.1 and B.2. Instead, for the weakly supervised case trained models are not available, so we reproduce the models as in [41] training them on Shapes3D and on other datasets that can be found in Appendix B.3.
@@
Fig. 1 (Left) shows the metric scores for different values of $\beta$ keeping the different FoV separated: the FoVs less affected by reconstruction (e.g. PosX and PosY) exhibit an increasing disentanglement score as $\beta$ grows. On the other hand, Shape and Orientation present a maximum value around $\beta=6$ and then decrease because they are more susceptible to the reconstruction quality, which degrades for larger values of $\beta$. In Fig. 1 (Center Left) an association matrix S is generated from one of the unsupervised models ($\beta$-VAE) trained with $\beta=6$: Shape and Orientation are encoded in the same dimension (overlapping) and produce lower values because of the reconstruction, while PosX and PosY are encoded in multiple dimensions and mostly overlapping. Scale does not seem to be well represented. In Fig. 1 (Center Right) we report the association matrix obtained by a weakly-supervised model (Ada-GVAE): the factors PosX, PosY and Scale are disentangled while Shape and Orientation are encoded in the representation with high intensity in different dimensions, with overlaps. Fig.1 (Right) shows OMES values ($\alpha=0.5$) computed over synthetic association matrices $S$, obtained by perturbing the ideal (diagonal) one. The perturbation aims to simulate 3 scenarios: noisy matrices where the disentanglement can be more or less strongly derived; models exhibiting partial disentanglement; models with no disentanglement. As it can be appreciated, the metric score nicely reflects the disentanglement intensity.
@@
Agreement with performance metrics. Similarly to [41], we consider the more informative weakly-supervised setting and discuss the rank correlation between our metric and performance evaluations (ELBO, reconstruction loss, and test error of FoVs classifier). Our analysis, reported in Fig. 2 (right) shows that OMES performs similarly to DCI, negatively correlated with the Reconstruction loss, and positively with the ELBO. It also correlates with the performances of GBT10000 (the classifier we will use in the experiments) while it mildly does with MLP10000. This empirical evidence is in line with what was observed in [11]. The correlation with the classification score is a sign that OMES is able to capture the property of expliciteness of the representation, although it is not directly measured by our metric. It is worth mentioning the correlation of OMES with the performance metrics is more stable than what is obtained by other metrics across different datasets (see the Appendix B.3).
@@ 3 Transferring disentangled representations
Fully unsupervised disentangled representation learning has been shown unsatisfactory in many scenarios [40]. However, annotating the FoVs can be a very critical and uncertain process. In this section, we propose a general-purpose methodology for transferring disentangled representations learned from supervised synthetic or simulated data to an unsupervised dataset (in terms of the FoVs). This approach allows us to evaluate the effectiveness of disentangled representations transfer, and its potential in real-world applications.
@@ 3.2 Datasets
Isaac3D [47] is a synthetic dataset of a 3D scene of a kitchen where a robot arm is holding objects in a variety of configurations. It is characterized by 9 real-world complex FoVs, including robot movements, camera height, environmental conditions (e.g. lighting).
@@
Coil is derived from Coil100 [46]. The original dataset contains 7200 real color images of 100 objects. The objects were placed on a motorized turntable against a black background. The turntable was rotated to vary object pose w.r.t. a fixed camera, producing self-occlusions and 2D silhouette changes. We augment the original dataset with two additional FoV, a planar rotation (9 angles) and a scaling (18 values). Therefore, we identify 4 FoV (Objects, Pose, Rotation and Scale) that, by construction, are independent. To consider in our analysis a real dataset visually related to dSprite, we derived a binary version of Coil, called Coil(bin), by applying Otsu’s thresholding [49].
@@ 3.3 Experimental analysis
The column Pruned highlights the two different representation modalities: if the classifier is trained on the whole representation (), or using only one dimension, i.e. the one showing the strongest encoding of a certain FoV according to the OMES metric (). As already mentioned, a good performance of the former is an indication of explicitness, while the latter is a positive sign of compactness. Tables also report metrics assessing Modularity (our MES and DCI) and Compactness (our OS and MIG).
Note that we exploit the interpretability of OMES in the transfer learning process to select the most representative dimension of the representation for the classification (the “Pruned” columns).
Table 2: Quantitative evaluation of transferred disentangled models using the dSprites family of datasets. We transfer from a Source (ST) to a Target Dataset (TD). We report the average classification accuracy obtained with GBT on the full and the pruned representations (see text). The last columns on the right report a comparison between disentanglement metrics, including MES and OS.
@@
(1) Synthetic to synthetic. As a baseline, we consider the case in which both Source and Target datasets are synthetic and we have access to the annotation of the FoVs, they are DRL-compliant. If Source and Target have the same FoVs (S=dSprites with T=Noisy-dSprites or S=Color-dSprites with T=Noisy-Color-dSprites, see Table 2) we observe that pruning the representation to just one dimension maintains, on average, stable performances. This shows that the compactness of the representation is preserved for the Target dataset, both before and after fine-tuning.
Fine-tuning allows for improved performance in terms of explicitness preserving the remaining properties of the representation, also in the case of the pruned representation. The Orientation FoV is difficult in these datasets as it suffers from reconstruction errors. We increase complexity by adding a new FoV to the Target dataset (S=dSprite with T=Color-dSprite, see Table 2).
All FoVs in common between Source and Target are effectively classified, again except Orientation. As for the new FoV (Color), we report lower performances, but we can appreciate a significant improvement with fine-tuning if we exploit a global representation. Instead, we observe a lower improvement with the pruned representation, suggesting that the new factor is not encoded in one single dimension.
@@
Table 3: Target dataset: Shapes3D (see Table 2).
@@
Isaac3D 62.8 (+26.2) 62.6 (+28.2) 39.8 (+37.9) 30.5 (+10.3) 51.9 (+35.2) 21.8 (+28.1) 44.9 (+27.6) 24.3 (+6.1) 3.6 (+16.7) 21.2 (+8.1) 1.5 (+5.2)
44.8 (+8.3) 42.5 (+14.4) 26.1 (+16.5) 22.3 (+6.6) 41.6 (+20.3) 13.8 (+14.5) 31.8 (+13.4)
We finally reason on the gap between Source and Target datasets in terms of complexity. When the Source is simpler than the Target but still they have some FoVs in common, possibly with different appearances, (e.g. S=Shapes3D, T=Isaac3D, Table 9 and Table 10) we can appreciate the effectiveness of transfer and fine-tuning for all metrics. Conversely, when the Source is much more complex than the Target (e.g. when S=Isaac3D, T=Shapes3D) one could expect the richness in the Source to be directly transferrable to the simpler Target. However, we observe that the finetuning is still beneficial for all the disentanglement metrics. This can be explained by the “domain” dependence of VAE models.
Discussion. Disentanglement transfers well between synthetic datasets with the same FoVs, w.r.t. all the properties. If the Target includes new FoVs, fine-tuning is necessary for the new FoV, but also for the entire representation, as compactness and modularity are partially degraded by the new FoV. When the Source and Target become significantly different, fine-tuning is also beneficial. We can conclude that when both Source and Target are synthetic and DRL-compliant, the properties of disentangled representation are preserved before and after fine-tuning, especially when the datasets have FoVs in common even though they have different appearance.
@@
With the same Target, we assess a representation which is incorporating some level of 3D information ( S=Shapes3D, with T=Coil, Table 4). This choice does not bring any benefit since the synthetic Shapes3D includes a very simplified form of pose variation. Orientation accuracy degrades, as it is not captured by the Source (indeed, with fine tuning this performance improves). Considering the Target dataset, it clearly presents several new challenges w.r.t. Source ones, we produce a last experiment with a simplified binary version of the Target, meant to be more similar to the binary images in dSprite. In this case, some FoVs are very well represented (Orientation and Scale), while Object presents low performances due to the decrease in the image descriptive power caused by binarization.
We now consider another real Target, RGBD-Object (see Table 5). As for the former, we consider Color-dSprite as a synthetic source: here, the same considerations about the Pose we discussed before are valid, and on the other FoVs, we observe the global representation is effective (and marginally improved by fine-tuning), while the pruned representation leads to lower performances, as a sign the representation is not perfectly disentangled on the Target. Here again, we investigate the possibility of transferring to simplified versions of the dataset (in this case, we consider the Depth channel only). The FoV Elevation improves significantly as it is not directly affected by color information. Concerning Modularity (Table 4 and Table 5), fine-tuning seems to have a small influence. In particular, we observe a small degradation for RGBD-Objects (Table 5), coherent with the challenges of a real dataset with unknown/hidden factors perturbing the encoding of the FoV.
Table 4: Target dataset: Coil100 and variants (see Table 2).
@@
Table 5: Target dataset: RGBD-Objects and variants (see Table 2).
@@
(3) Real to Real. We conclude by discussing the possibility of transferring from a DRL-compliant real dataset to another real one. As a first task, we consider as a Source a simplified version of the Target (specifically, S=Coil-binary, with T=Coil): the source should encode the factors not related to RGB, while the finetuning should improve the disentanglement and the explicitness of the representation. However, this is not the case with Coil100, whose representations degrade the Modularity, and the finetuning only affects the entire representation.
We then consider a larger variation between Real Source and Real Target (specifically, S=Coil with T=RGBD-Object, see Table 5): we obtain similar results to those of Color-dSprites as Source Dataset (comparable Explicitness), with a reduction on the performances obtained by the pruned representation. Notice that adopting the binary Coil as a Source causes only a limited reduction in Explicitness, and this was somewhat unexpected as we have a large gap in complexity between Source and Target. Our experiments did not consider RGBD-Objects acting as a Source dataset, not being DRL-compliant.
@@
4 Limitations
A limitation of our current work is the adoption of a specific family of approaches (VAE-based). The generalization of our finding to more recent vector-based approaches (e.g. [71, 62, 48, 36]) needs further investigation. However, each family of approaches for disentanglement learning follows specific paradigms that may require tailored designs for transfer learning. In other words, while the general transfer methodology is still applicable, it might need proper tuning to perform optimally depending on the particular learning approach.